# Exploration of Masked and Causal Language Modelling for Text Generation

### v1

@@ IV-C Downstream Evaluation Tasks
In this section, we evaluate the generated texts used in three downstream tasks: NER, Text Classification, and Authorship Verification. These results will suggest the relationship between the quality of the generations and their usefulness in a practical setting, assessing the potential for use as datasets for different NLP tasks. In Section III-D we explained how each downstream task was designed.
@@ IV-C1 Downstream NER Task
Figure 11 shows that while the models trained on the real data perform slightly better (orange line), the models trained on the generated data also perform well. The graph shows a similar behaviour as the quantitative metric graphs. It shows good scores for high ratios, and it descends, although less decisively, for lower contexts. In contrast with the quantitative metrics, however, synthetic data generated via stopwords and punctuation as context can train the models to a comparable performance to random masking. For MLM models, NER tokens provide enough context to train the model effectively.
What is remarkable about these results is that even when the ratio is at 100%, the f-1 sharply decreased. However, it’s still able to effectively train the model, achieving an f-1 of almost 75%. This contrasts with the low quantitative scores achieved by this ratio. This shows that to train NER models effectively, the quality of the generated samples has a limited impact on the performance.
@@ IV-C2 Downstream Text Classification Task
Figure 12: Downstream Text Classification Evaluation Results.
The results show that the models trained on the generated data perform significantly differently based on the quality of the generated texts. The worst model had an f-1 score of 0.42, and the worst model was 0.36. Even more evidently than in the NER task, the quality of the text does not correlate with the performance of the model; as a matter of fact, the Roberta-large 100% masking ratio performs better than most other ratios and punctuation for t5-large performs even better than the original training data.
It should be noted that given the generation constraints, we could train the model only on 200 samples; therefore, the models don’t perform excellently in this task.
IV-C3 Downstream Author Verification Evaluation
Figure 13: Downstream Authorship Verification Evaluation Results.
The generations show performances comparable to the quantitative metrics for 3 out of the 4 models. This is a good sign which means that the models can leverage the context appropriately. For the MLM models, however, the performance decreases sharply at the lower end of the ratios (90 - 100%). T5-large, in contrast, shows no correlation between the number of matches and the amount of context given, meaning that the model is indifferent to the amount of context given.
What is peculiar is also the fact that punctuation alone obtains the highest performance of all for the CLM models (T5 & Bart); however, when looking at the actual generations, we realized that the models show clear hallucinations when given just punctuation, in fact all the generations are just ",,,,," leading the prediction model to predict correctly predict "Yes" (text 1 and the reconstructed text come from the same author) but with low confidence.
With regard to stopwords and stopwords&punctuation, they show fairly good performance, placing in the middle of the graph. Punctation does not give enough context to the MLM models and shows very low performance, which is shown as the lowest in the Figure.
@@ IV-D Discussion
The quantitative and qualitative evaluations of this study, as well as the performance shown in the downstream tasks, provide valuable insights into the performance of the different language models and which are the best masking approaches for text generation.
Our main finding is that 1) MLM consistently outperformed CLM in text generation across all three datasets. The quantitative metrics indeed showed that the texts generated by MLM models were more aligned with the original reference texts. The qualitative human evaluation showed that MLM generations had better coherence and were more grammatically correct.
2) Incorporating domain-specific knowledge in the generation models did not improve performance. This was shown by the scores obtained with BiomedNLP-PubMedBERT and SciFive-Large-Pubmed_PMC, two models focused on the biomedical domain, did not achieved better scores than the general models in any of the quantitative metrics.
3) We found no strong correlation between the quality of the generated texts and the performance in the downstream tasks. While most models trained on the generations of higher quality did perform well in the downstream tasks, some of the models trained on lower-quality data (e.g., 100% masking ratio) still achieved good results. We believe that this is a very interesting discovery, and it suggests that downstream tasks are more resilient to the noise that the generative models add with their generations. It is also possible that the synthetic generations make the models more robust since they introduce significant variability in the training sets.
4) As expected, random masking with lower masking ratios (preserving more context) led to higher-quality generations. However, even with limited context, such as NER or stopwords_punctuation masking, MLM models could generate texts of good overall quality.
@@ V Conclusions and Future Work
In this paper, we compared CLM and MLM text generation across medical, movie synopsis, and authorship verification datasets. We explored different masking approaches to determine the most effective ones. We then assessed the quality of the generations first with quantitative and qualitative evaluations, and then we assessed their usability by running three downstream tasks. Our results established that MLM can generate better text than CLM.
@@ V-A Limitations
Our findings provide valuable insights into the performance of MLM and CLM approaches for text generation. However, in this section, we want to acknowledge some of the limitations that may impact the generalisability of our results.
Computational Efficiency: The iterative process of generating text using MLM models, particularly with high masking ratios, can be computationally expensive compared to autoregressive CLM models. Each iteration requires feeding the entire masked sequence through the model, which can be time-consuming for longer texts. This is because The current libraries are not optimized for generating texts using MLM models.
Quantitative Evaluation Metrics: The quantitative evaluation relied upon n-gram-based metrics because they provide a valuable indication of text quality; however, it should be noted that they were initially designed for tasks such as machine translation and text summarisation. As such, these measures may not fully align with the kind of text that we are looking for. In addition, it should be the MLM models had an advantage in this evaluation because part of the real tokens were already in the right place. Therefore, it is possible that a higher score does not correspond to better text quality in this case.
Length of the Texts: The BERT family of models, including those we employed, has a maximum input sequence length of 512. This may limit the application of MLM-based text generation in domains that usually involve longer texts.
Size of the Models: Due to computational resources and the fact that encoder-only models are generally small, we had to restrict our study to models that have hundreds of millions of parameters. Therefore, their performance may be limited compared to larger state-of-the-art models with billions of parameters, and further research should focus on that instead.
@@ V-B Future Work
Creative Writing: Future work could involve our method to generate more unstructured text generated in tasks such as creative writing. This will help us understand whether our approach can not only enhance the quality of the generations but also their creativity.
Iterative Refinement: Iterative refinement [32] involves gradually improving the initial generations over multiple iterations until good text quality is achieved. When applied to MLM, this process could entail identifying and masking problematic tokens or text sequences and then regenerating those parts to improve the quality of the final output.
More Downstream Tasks: This paper showed that texts of higher quality do not necessarily lead to higher performance in the downstream tasks. In this work, we focused on NER, text classification and authorship verification; however, future work should include a broader range of NLP tasks, such as sentiment analysis, question answering and POS tagging. By evaluating the performance of models trained on MLM-generated data across a diverse set of tasks, we will be able to show more definitively the versatility and robustness of this method.
On Medical Data: MLM models showed the best performance compared to CLM in the medical dataset so future research could investigate the application of MLM for data augmentation and synthetic data generation in the biomedical domain further. This is, in fact, the domain that needs this kind of data the most because it is generally hard to access due to privacy concerns [7]. Future research could involve more masking strategies and text types, as well as evaluating the generated data on a range of medical downstream tasks. Another critical aspect to add to this research would be to collaborate with actual clinicians to assess the plausibility of the generated text along with its quality; this could provide us valuable insights into the possible use of this approach in clinical settings.

### v2

@@ IV-C Downstream Evaluation Tasks
In this section, we evaluate the generated texts by utilizing them in three downstream tasks, namely NER, Text Classification, and Authorship Verification. These results will suggest the relationship between the quality of the generations and their usefulness in a practical setting, assessing the potential to use them as synthetic datasets for different NLP tasks. The design of each downstream task was explained in Section III-D.
@@ IV-C1 Downstream NER Task
Figure 13 demonstrates that, although models trained on real data perform slightly better (orange line), models trained on generated data could perform decently. The graph exhibits a similar trend to the quantitative metric graphs, showing high scores for higher ratios and a less decisive decline with reduced contexts. Unlike the quantitative metrics, synthetic data generated using stopwords and punctuation as context could train the models to achieve performance comparable to random masking. For MLM models, NER tokens could provide sufficient context to train the model effectively.
Remarkably, even when the masking ratio reaches 100%, the F1 score, though sharply decreased, remains effective enough to train the model, achieving an F1 score of nearly 0.75. This contrasts with the low quantitative scores observed at this ratio. Therefore, these results suggest that for training NER models, the quality of generated samples has a limited impact on performance.
Figure 14: Downstream Text Classification Evaluation Results.
Figure 15: Downstream Authorship Verification Evaluation Results.
@@ IV-C2 Downstream Text Classification Task
The results in Figure 15 indicate that the models trained on generated data exhibit significant performance variation based on the quality of the generated texts. The best model achieved an F1 score of 0.42, while the worst scored 0.36. Unlike in the NER task, the quality of the text does not strongly correlate with model performance. In fact, the RoBERTa-large model with a 100% masking ratio outperforms most other ratios, and the T5-large model with punctuation masking performs even better than when trained on the original data.
It is important to note that, due to hardware constraints, we could only train the models on 200 samples, which limits their overall performance in this task.
IV-C3 Downstream Author Verification Task
As we can observe from Figure 15, the generated texts exhibit performances comparable to the quantitative metrics for three out of the four models. This indicates that the models can leverage the context appropriately. However, for the MLM models, performance decreases sharply at the higher masking ratios (90-100%). In contrast, T5-large shows no correlation between the number of matches and the amount of context provided.
Interestingly, punctuation alone achieves the highest performance for the CLM models. However, upon examining the actual generations, we noted that the models exhibit clear hallucinations when given just punctuation. Specifically, all generations are merely ",,,,,", leading the prediction model to correctly predict "Yes" with low confidence.
Regarding stopwords and the combination of stopwords and punctuation, they demonstrate fairly good performance, positioning them in the middle of the graph. Punctuation alone does not provide sufficient context for the MLM models, resulting in very low performance, as depicted at the bottom of the figure.
@@ IV-D Discussion
The quantitative and qualitative evaluations of this study, as well as the performance observed in the downstream tasks, provide valuable insights into the effectiveness of different language models and suggest optimal masking approaches for text generation. Our main findings are given below:
(1) MLM Outperforms CLM: MLM consistently outperformed CLM in text generation across all three datasets. Quantitative metrics showed that texts generated by MLM models were more aligned with the original reference texts. In addition, qualitative human evaluation indicated that MLM generations had better coherence and grammatical accuracy.
(2) Domain-Specific Knowledge does not Improve Performance: Incorporating domain-specific knowledge in the generation models did not enhance performance. This is evidenced by the scores from BiomedNLP-PubMedBERT and SciFive-Large-Pubmed_PMC, two models focused on the biomedical domain, which did not achieve better scores than the general models at any quantitative metrics.
(3) There is no strong correlation between the quality of generated texts and performance in downstream tasks. While most models trained on higher-quality generations performed well, some models trained on lower-quality data (e.g., 100% masking ratio) also achieved decent results. This finding suggests that downstream tasks are more resilient to noise introduced by generative models. It is also possible that synthetic generations enhance model robustness by introducing significant variability in the training sets.
(4) Effectiveness of Random Masking with Lower Ratios: As expected, random masking with lower masking ratios (preserving more context) resulted in higher-quality generations. However, even with limited context, such as NER or stopwords_punctuation masking, MLM models could still generate texts of good overall quality.
@@ V Conclusions and Future Work
In this paper, we compared CLM and MLM for text generation across medical, movie synopsis, and authorship verification datasets. We explored different masking approaches to determine the most effective ones. We then assessed the quality of the generations; first through quantitative and qualitative evaluations, and then by assessing their usability on three downstream tasks. Our results established that MLM models could generate better texts than CLM models.
@@ V-A Limitations
Our findings provide valuable insights into the performance of MLM and CLM approaches for text generation. However, in this section, we want to acknowledge some of the limitations that may have impacted the generalisability of our results.
Quantitative Evaluation Metrics: The quantitative evaluation relied upon n-gram-based metrics because they provide a valuable indication of text quality. However, it should be noted that they were initially designed for tasks such as machine translation and text summarization. As such, these measures may not fully align with our kind of texts. In addition, it could be that MLM models had an advantage in this evaluation as part of the true tokens were already provided to the models.
Length of the Texts: The BERT family of models, including those we employed, have a maximum input sequence length of 512 tokens. This may limit the application of MLM-based text generation in domains that usually involve longer texts.
Size of the Models: Due to computational resources and the fact that encoder-only models are generally smaller, we had to restrict our study to models that have hundreds of millions of parameters. Therefore, their performance may be limited compared to larger state-of-the-art models with billions of parameters. Further research should focus on that instead.
@@ V-B Future Work
Creative Writing: Future work could involve our method to generate more unstructured texts in tasks such as creative writing. This would demonstrate whether our approach could also enhance generations’ creativity, on top of their quality.
Iterative Refinement: Iterative refinement [31] involves gradually improving the initial generations over multiple iterations until good text quality is achieved. When applied to MLM, this process could entail identifying and masking problematic tokens or text sequences, and then re-generating those parts to improve the quality of the final output.
More Downstream Tasks: This paper showed that texts of higher quality do not necessarily lead to higher performance in the studied downstream tasks. As we focused on NER, text classification and authorship verification; future work should include a broader range of NLP tasks, such as sentiment analysis, question answering and POS tagging. By evaluating the performance of models trained on MLM-generated data across a diverse set of tasks, we will be able to show the versatility and robustness of this method with more certainty.
On Medical Data: MLM models showed the best performance compared to CLM in the medical dataset. Therefore, future research could further investigate the application of MLM for data augmentation and synthetic data generation in the biomedical domain. This domain particularly benefits from such approaches due to the challenges of accessing data because of privacy concerns [7]. Future research could involve additional types of texts and masking strategies, as well as evaluating the generated data on a range of medical downstream tasks. Another critical aspect would be to collaborate with clinicians to assess the plausibility of the generated texts along with their quality. This could provide us valuable insights into the possible use of this approach in clinical settings.
Acknowledgements
We thank Prof Juntao Li from Soochow University’s NLP group for the valuable feedback on our work.
author contributions
NM: experiments and writing (the original thesis, which this article is based upon); SB: writing; LH: article structure and writing (extracting from the thesis); GN: supervision.